I need to write an equation to this problem
The perimeter of a triangle is 51 cm. The longest side is twice the length of the shortest side. The third side is 3 com longer than the shortest side. What is the length to each side?

Answer by stanbon(75887)

The perimeter of a triangle is 51 cm. The longest side is twice the length of the shortest side. The third side is 3 cm longer than the shortest side. What is the length to each side?
-----------
Let shortest side be "x".
Then longest side is "2x".
And 3rd side is "x+3".
EQUATION:
x+2x+x+3=51
4x+3=51
4x=48
x=12 cm (length of shortest side)
2x=24 cm (length of longest side)
x+3=15 cm (length of 3rd side)
Cheers,
Stan H.